OpenStudy (anonymous):

solve for x: 4√5x + 1 = 2

OpenStudy (anonymous):

1/80

OpenStudy (anonymous):

well move the 1 to the other side 4sqrt(5x)=1 then divid by 4 so sqrt(5)=1/4 then take the square of both sides to get rid of the radical 5x=1/16 then divid by 5 on both sides to get x alone

OpenStudy (anonymous):

so x = (1/16)(1/5) =1/80

OpenStudy (anonymous):

YEAH BRO